OpenStudy (anonymous):

tan(theta)=−3/4 and cos(theta)>0. Find sin(theta), cos(theta), cot(theta), csc(theta), and sec(theta).

OpenStudy (anonymous):



OpenStudy (anonymous):

find the hypotenuse by pythagoras or memory then take the appropriate ratios, being mindful of what quadrant you are in so you can put the right sign \(+\) or \(-1\) in your answer

OpenStudy (anonymous):

It says cos(theta)>0 So..it must be in 4th quadrant

OpenStudy (anonymous):

That's what I don't understand. How do I know which quadrant I am in when cos theta is > 0?

OpenStudy (anonymous):

OH. Okay. Let me try :)

OpenStudy (anonymous):

forget about the signs until the answer sides of a triangle are positive you are in quadrant 4 because cosine in positive but sine is negative.

OpenStudy (anonymous):

from the picture we see cosine is \(\frac{3}{5}\) and you know it is positive because you are told

OpenStudy (anonymous):

oops i mean cosine is \(\frac{4}{5}\) sorry

OpenStudy (anonymous):

sine is \(-\frac{3}{5}\) you get the \(\frac{3}{5}\) from the triangle, and it is negative because you are in quadrant 4 etc

OpenStudy (anonymous):

OH i get it okay. one second. THANK YOU!

OpenStudy (anonymous):

so sin = -3/5; cos = 4/5; cot 4/3; csc -4/5; sec 5/4?

OpenStudy (anonymous):

cot = -4/3**

OpenStudy (anonymous):

Yup....

OpenStudy (anonymous):

your cosecant is wrong, all others are right

OpenStudy (anonymous):

Sec and Cos are only +ve..

myininaya (myininaya):

and the cotangent

myininaya (myininaya):

oh you corrected it

OpenStudy (anonymous):

cosec = 1/ sin x

OpenStudy (anonymous):

cosec x = -5/3

OpenStudy (anonymous):

I got: sin -3/5; cos 4/5; cot -4/3; csc -5/3; sec 5/4 correct?

OpenStudy (anonymous):

thank you!